OpenStudy (anonymous):

Solve x² – 121 = 0? There is no solution. –11 11 11, –11

OpenStudy (kinggeorge):

Notice that you can factor any binomial of the form \(x^2-a^2\) as \((x-a)(x+a)\). Since \(121=11^2\), you know that \(x^2-121=(x-11)(x+11)=0\). So it has two solutions, \[x=11\]\[x=-11\]